A submersible is spherical with a diameter of 4.00
Answer #1

a) Buoyant force =1000*9.8*(4/3*3.14*4^3)

=2.626*10^6 N

b) Minimum mass required =1000*(4/3*3.14*4^3)

=2.6795*10^5 Kg